The 2.5-kg collar travels with negligible friction
Answer #1

from the Newton second law of motion, the accleration of the collar is

a = F/m = 59 N/2.5 kg =23.6 m/s^2

from the kinmatic equation

vB^2- uB^2 = 2aR

since u = 0

vB = root 2aR

= root 2 ( 23.6 m/s^2) (2.8 m)

=11.49 m/s